# quality of service 5G fifth generation

### v1

@@ IV Methodology
Our primary focus is to address resource allocation problem within O-RAN architecture, and we propose a novel approach based on semi-supervised learning. Our objective is to optimize the sum of rates among UEs while considering constraints, including transmission power and PRB allocation. To accomplish this, we compile a dataset comprising diverse resource allocation scenarios, encompassing various combinations of objective functions and constraints. We generate this dataset using an exhaustive search algorithm (ESA), which explores the entire feasible solution space.
Our research primarily delves into the resource allocation problem within the O-RAN architecture, and we introduce an innovative approach based on semi-supervised learning. Our primary goal is to enhance the cumulative rates for UEs while adhering to specific constraints such as transmission power, delay, and PRB allocation. To achieve this, we curate a comprehensive dataset that encompasses a wide range of resource allocation scenarios, covering diverse combinations of objective functions and constraints. This dataset is generated through the implementation of an ESA, which thoroughly explores the entire feasible solution space.
In our approach to addressing the resource allocation problem, we employ a deep neural network (DNN) model integrated with a contrastive loss function. This DNN model efficiently manages the complexities of resource allocation and acquires the ability to map input features to resource allocation decisions. Through the training of our model on the dataset we’ve created, we fine-tune it to estimate resource allocation decisions that fulfill the requirements of the optimization problem. This semi-supervised learning algorithm, which combines a VAE with a contrastive loss, is denoted as ”SS-VAE.”
We aim to evaluate the performance of our learning-based approach using multiple regression-based performance metrics to assess its robustness. Additionally, we will validate the effectiveness of our model by comparing its outcomes with those of the ESA in different scenarios. Moreover, while the ESA is the optimal algorithm for resource allocation decisions, its complexity makes it impractical for deployment in real-world environments. Therefore, we will also compare the performance of our proposed algorithm with state-of-the-art methods known for their practical deployment capabilities, such as DRL. This comprehensive assessment will demonstrate the method’s reliability across diverse scenarios.
Figure 2 illustrates the overview of our proposed methodology for training the SS-VAE and comparing its results with those of the benchmark algorithms.
Figure 2: The overview of proposed methodology.
@@ IV-A Exhaustive Search Algorithm
The ESA is a well-established optimization approach for solving resource allocation problems in wireless networks. Its fundamental concept involves generating all possible resource allocation combinations and selecting the best one based on a predefined objective function [33]. ESA serves as a valuable benchmark solution, providing a ground truth against which results from other methods can be validated and compared. Its advantages include simplicity of understanding and implementation, along with the assurance of yielding an optimal solution within a finite time frame.
Nonetheless, ESA’s primary limitation lies in its computational complexity, especially when addressing resource allocation problems with extensive dimensions and intricate constraints. The computational burden grows exponentially with the size of the problem, rendering ESA impractical for large-scale resource allocation tasks [34]. Despite this drawback, ESA remains a popular choice as a baseline for comparing various optimization algorithms and offers a crucial reference for evaluating new optimization strategies.
To assess the performance, validity, and robustness of the proposed learning-based method, we will compare its results with those generated by ESA. Additionally, a subset of solutions for the resource allocation and network slicing optimization problem within O-RAN will be generated using ESA. This labeled dataset will be utilized to train the supervised learning component of the proposed method.
@@ IV-B Deep Reinforcement learning Algorithm
As discussed in Section II, DRL algorithms have proven to be effective solutions for addressing resource allocation problems. These algorithms have been successfully deployed in various real-world wireless network environments, demonstrating their practicality and their ability to significantly improve both technical and business metrics [35]. Consequently, we utilize DRL, where an agent interacts with the environment to make resource allocation decisions and receive related rewards. These rewards are tied to the objective function outlined in Problem 17, considering its objectives and constraints. Next, we model the problem as a Markov Decision Process (MDP) and introduce the DRL algorithm to address the resource allocation challenge in network slicing. This approach serves as an extra means to evaluate the robustness and our primary proposed solution.
IV-B1 Markov Decision Process
In the specified resource allocation problem (Equation 17), an O-RAN orchestrator is responsible for determining PRB allocation decisions and determine the corresponding RU for each UE for entw. This problem can be represented as a MDP denoted by $(\mathfrak{S},\mathfrak{A},\mathfrak{P},\mathfrak{R})$, where $\mathfrak{S}$ represents the state space, $\mathfrak{A}$ is the action space, $\mathfrak{P}(\mathfrak{s}^{\prime}|\mathfrak{s},\mathfrak{a})$ signifies the transition probability from state $\mathfrak{s}$ to state $\mathfrak{s^{\prime}}$ by taking action $\mathfrak{a}$, and $\mathfrak{P}$ represents the reward function [35, 36, 37, 38], which are elaborated for our specific problem.
@@ IV-B1 Markov Decision Process
    State: The state at time $t$ ( $\mathfrak{s}(t)\in\mathfrak{S}$), is represented by $\{\mathfrak{s}_{u}(t)\}_{t=1}^{N}$, where $\mathfrak{s}_{u}(t)$ indicates the state of UE $u$ ($u=1:U$). Here, $\mathfrak{s}_{u}(t)$ is a binary indicator: it equals 1 if the data rate requirement of UE $u$ is met and 0 otherwise. It’s important to note that each data rate corresponds to a specific transmission power level.
@@
    Reward: The reward function serves as a pivotal element in assessing the gains or expenses associated with a given state-action pair. It provides guidance to the agent’s decision-making process, indicating the appeal of different actions across various states. In alignment with earlier studies [35, 36, 37], our method characterizes the reward function as a combination of the objective function and the constraints relevant to the considered problem (17). Therefore, we establish the reward function for our specific problem as follows:
@@
    where, $\Theta_{r}$, $\Theta_{const.}$, and $\Theta_{bias}$ represent the respective weights assigned to the objective function (Data Rate), constraints, and the bias value.
The primary objective of the agent is to determine the optimal policy, denoted as $\pi:\mathfrak{S}\rightarrow\mathfrak{P(A)}$, for the MDP. This policy maps any state $\mathfrak{s}\in\mathfrak{S}$ to a probability distribution $\pi(\cdot|\mathfrak{s})$ over the set of possible actions $\mathfrak{A}$. The goal is to find a policy that maximizes the cumulative reward obtained through dynamic learning from acquired data. The policy $\pi$ can be influenced and controlled by decision makers.
The corresponding value function $V(\pi):\mathfrak{S}\rightarrow\mathfrak{R}$ is defined as the cumulative discounted reward achieved by following the actions prescribed by the policy $\pi$ when starting from a specific initial state. Additionally, the action value function $Q^{\pi}(\mathfrak{s},\mathfrak{a})$ is introduced to represent the expected cumulative reward when taking action $a$ in state $\mathfrak{s}$, considering that the policy $\pi$ is being followed. This can be expressed as follows:
$V^{\pi}(\mathfrak{s})=\mathbb{E}[Q^{\pi}(\mathfrak{s},\mathfrak{A})|\mathfrak{A}\sim\pi(.|\mathfrak{s})],$ (30)
$Q^{\pi}(\mathfrak{s},\mathfrak{a})$ $=\mathbb{E}[\sum_{t=0}^{\inf}\gamma^{t}\cdot\mathfrak{R_{t}}|\mathfrak{S_{0}}=\mathfrak{s},\mathfrak{A_{0}}=\mathfrak{a}]$
$=\mathfrak{r}(\mathfrak{s},\mathfrak{a})+\gamma\;\mathbb{E}[V^{\pi}(\mathfrak{S}^{\prime})|\mathfrak{S}^{\prime}\sim\mathfrak{P}(.|\mathfrak{s,a})],$ (31)
where, $\mathfrak{R}_{t}$ represents the immediate reward obtained at time $t$, and $\gamma$ (where $10<\gamma<$ 1) denotes the discount factor applied in the computation of future rewards. The discount factor determines the weight of future rewards in the decision-making process.
@@
Exploring the dynamics of the environment, especially when accounting for variable channel conditions and calculating transition probabilities, is a highly complex task. Determining the value function in such scenarios presents a significant challenge. To address MDPs, two primary strategies are employed: model-based and model-free approaches. In instances where there is uncertainty surrounding transition probabilities and reward functions, the model-free approach of Q-learning proves invaluable for solving the MDP [38].
IV-B2 Deep-Q-Network Algorithm
In our considered problem, we face the challenge of managing the states space ($\mathfrak{S}$) and actions space ($\mathfrak{A}$) that are inherently high-dimensional and complex. This complexity arises from continuous variable like transmission power, as well as the substantial number of UEs and PRB resources. Traditional Q-learning methods struggle in such scenarios, as many state-action pairs are rarely encountered, leading to inefficient learning processes [35, 38].
To overcome this challenge, we turn to the Deep Q-Network (DQN) approach to tackle our resource allocation problem and optimize parameters such as transmission power, alpha, and beta. DQN offers distinct advantages in handling these complex, high-dimensional state-action spaces by utilizing a deep neural network to approximate the $Q$-function. DQN’s capacity to generalize from prior experiences and its use of techniques such as experience replay and target networks make it an apt choice for addressing resource allocation challenges within our problem domain.
DQN employs experience replay to store into replay memory ($\mathbb{M}$) and sample transitions from the memory ($\mathfrak{S_{t}},\mathfrak{A_{t}},\mathfrak{R_{t}},\mathfrak{S_{t+1}}$) to train the neural network with stochastic gradient descent. This method helps acquire uncorrelated samples from a temporally correlated MDP trajectory, enhancing gradient estimation accuracy. A target network $Q_{\theta^{*}}$ is also used, where independent samples are obtained from the replay memory. These samples are used to update the $Q$-network’s parameter, improving learning stability and efficiency [38]. We compute the target value and update $\theta$ by the gradient of equations 32 33.
$Y=\mathfrak{R(s,a)}+\gamma\;\max_{\mathfrak{a}^{\prime}\in\mathfrak{A}}Q_{\theta^{*}}(\mathfrak{s}^{\prime},\mathfrak{a}^{\prime}),$ (32)
$L(\theta)=||Y-Q_{\theta}(\mathfrak{s},\mathfrak{a}||^{2},$ (33)
The parameter $\theta^{*}$ is updated every $T_{target}$ steps by setting it equal to $\theta$. The target network remains unchanged for $T_{target}$ steps and is then updated using the current weights of the Q-network.
@@ IV-B2 Deep-Q-Network Algorithm
It’s worth noting that the complete training algorithm for DQN is detailed in [38], and for our specific problem, we have utilized Algorithm 3 for training the DQN model.
IV-C Proposed Method
In recent times, DL techniques have found application in enhancing the accuracy and efficiency of resource allocation and transmission power control in wireless communication systems, as demonstrated in several studies [39, 40, 41, 42]. These approaches excel in learning complex, non-linear relationships between input features like channel matrices and output decisions, such as power and resource allocation. What distinguishes DL-based resource allocation is its ability to effectively manage the dynamic and intricate nature of wireless communication systems. These approaches can adapt to varying channel conditions and UE, resulting in improved network performance compared to traditional optimization methods.
In our proposed DL-based resource allocation method, we feed the channel matrix into a DNN. This network is responsible for learning optimal resource allocation decisions that adhere to network constraints while maximizing the defined objective function (weighted aggregated data rate). We also employ a flattened vector with dimensions $UB$, derived from the channel gain matrix, as input for a machine-learning model, which includes a VAE. This autoencoder utilizes the channel matrix to generate the power vector. The key innovation in our approach is the introduction of contrastive loss as an additional term in the VAE’s loss function. This addition refines the learning of representations, significantly enhancing the accuracy of solutions to resource allocation problems. Subsequent sections will provide a detailed exploration of each component in our proposed method.
@@ IV-C1 Variational Autoencoder
The VAE is a specific type of generative model frequently applied in unsupervised learning tasks. It functions as a DNN designed to generate new data samples resembling a provided set of training data. Utilizing a VAE confers a notable advantage: it adeptly captures the intrinsic distribution of the channel gain matrix. This acquired knowledge proves immensely valuable when confronting the resource allocation problem [43].
In a VAE, let’s denote the input data as $x$, a latent variable representing a lower-dimensional data representation as $z$, and $p_{\theta}(x)$ as a generative model representing the data distribution based on the latent variable. The primary goal of a VAE is to maximize the likelihood of the data given the model parameters, denoted as $\theta$. This likelihood can be mathematically expressed as:
@@
where, $q_{\phi}(z|x)$ corresponds to the encoder network, which models the posterior distribution of the latent variable given the input data. The term $D_{KL}(q_{\phi}(z|x)|p(z))$ represents the Kullback-Leibler (KL) divergence between this posterior distribution and the prior distribution of the latent variable.
The objective function can be enhanced by adjusting the parameters of both the generative model, denoted as $\theta$, and the encoder network, represented as $\phi$. This optimization process involves iteratively updating these parameters and computing the gradient of the objective concerning them, as outlined in [43]. Figure 3 illustrates the VAE architecture tailored for addressing the resource allocation and network slicing problem.
Figure 3: The architecture of proposed VAE.
To provide a more detailed explanation, a VAE comprises two primary components: an encoder and a decoder. The encoder accepts the channel gain matrix as input and transforms it into a lower-dimensional representation known as a latent code. Subsequently, the decoder takes this latent code as input and produces an estimation of the transmission power. These two components work in tandem during training to minimize a reconstruction loss. This loss metric quantifies the disparity between the actual transmission power and the predicted transmission power, guiding the network to generate more accurate power predictions.
Using a VAE for resource allocation offers several distinct advantages. Firstly, VAEs exhibit remarkable flexibility, enabling them to model intricate distributions. This characteristic is particularly advantageous for capturing the complex and non-linear relationship between channel gains and transmission powers. Secondly, VAEs excel at discerning the underlying structure within data, enabling them to make informed predictions even when confronted with previously unseen data [44]. Thirdly, the latent code generated by the encoder serves as a compressed representation of channel gains, which can be leveraged for diverse tasks like visualization, clustering, and anomaly detection.
Moreover, resource allocation problems are inherently non-convex and can be exceptionally challenging when using traditional optimization approaches. The introduction of VAEs provides a more versatile and effective means of tackling these intricate problems.
@@
Our dataset comprises a total of $Q$ training samples, divided into two categories: $Q_{L}$ labeled samples and $Q_{U}$ unlabeled samples. These samples are represented by pairs of channel matrices and their corresponding power allocation vectors, denoted as $H_{i}$ and $\Gamma_{i}$ for labeled samples (where $i$ ranges from 1 to $Q_{L}$), and $H_{j}$ for unlabeled samples (where $j$ ranges from 1 to $Q_{U}$).
It is important to note that the process of generating these input channel matrices is computationally less intensive compared to traditional methods, as documented in previous research [45].
@@
To train the supervised learning component of our algorithm, we employ the ADAM optimizer [46]. This optimizer facilitates the end-to-end training of the VAE using the $Q_{L}$ labeled samples. The loss function for the regression problem with ADAM optimization is formulated using the mean squared error (MSE) loss function, defined as follows:
$L_{supervised}=\frac{1}{\beta_{L}}\sum_{i\in\beta_{L}}||VAE(H_{i})-\Gamma_{i}||^{2}.$ (35)
In our resource allocation challenge, the correlation between the xApps dealing with power estimation and the PRBs allocation emerges from their shared foundation within the VAE framework. The VAE operates as a unifying space where both xApps converge, enabling a synergistic exploration and consideration of the pertinent features relevant to both power transmission and PRB allocations. Through this shared space, the xApps gain access to a common latent representation of the channel gain matrices. This shared understanding facilitates a cohesive analysis where the interdependencies between power estimation and PRB allocations are thoroughly examined, allowing for the simultaneous consideration of the critical features that impact both domains. This integration of information aids in making more comprehensive and informed resource allocation decisions, as it ensures that both xApps leverage a unified understanding of the underlying data to optimize their respective objectives. This coordination is pivotal, as it allows for a coordinated and synchronized approach to resource allocation within the O-RAN architecture.

### v2

@@ III System Model and Problem Formulation
This formulation can be seen as a well-known NP-hard problem, a multi-dimensional knapsack problem. Since we can reduce it to our resource allocation problem in polynomial time, it follows that it is also NP-hard. Given the NP-hard nature of the resource allocation problem, finding an exact solution in polynomial time is computationally infeasible for large-scale networks.
The system model is constrained by limited resources such as UE or RU and slice power and fronthaul capacity, preventing the aggregate throughput from exceeding its optimal value. This ensures that the objective function, representing aggregate throughput, has an upper bound and cannot increase infinitely. As a result, the algorithm used to solve the system will converge to an optimal solution. If the objective function is a strictly ascending function concerning the number of iterations, it will converge to its global optimum. The algorithm will converge to a local optimum if it is a non-monotonically ascending function. This guarantees the stability and convergence of the system model within the feasible region.
Notation Definition
$b,\mathcal{B},B$ Index, set, and the number of RUs
$u,\mathcal{U},U$ Index, set, and the number of UEs
$m,\mathcal{M},M$ Index, set, and the number of PRBs per RU
$s,\mathcal{S},S$ Index, set, and the number of slices
$h,H$ Index, and channel gain matrix between RUs and UEs
$\alpha_{u,b,s}$ UE $u$ association to RU $b$ in slice $s$
$\beta_{u,b,m,s}$ UE $u$ association to PRB $m$ of RU $b$ in slice $s$
$p_{u,b,m,s}$ Transmission power of PRB $m$ of RU $b$ in slice $s$
$R_{u,b,m,s}$ Rate achieved for UE $u$ in PRB $m$ of RU $b$ in slice $s$
$R_{u,s}$ Rate achieved for UE $u$ in slice $s$
$R_{u}$ Rate achieved for UE $u$
$w_{s}$ Priority weights for each service type
$P_{b}$ Transmitted power RU in the fronthaul link
$C_{b}$ Data rate in the fronthaul link
$\sigma_{1}^{2}$ Quantization Noise
$D^{pro}$ Overall propagation delay
$D^{fr,p}$ Propagation delay in the fronthaul link
$D^{mid,p}$ Propagation delay in the midhaul link
$D^{b,p}$ Propagation delay in the backhaul link
$L$ Length of the fiber link
$C$ Propagation speed of the medium
$D^{tr}$ Total transmission delay
$D^{fr,t}$ Transmission delay in the fronthaul link
$D^{mid,t}$ Transmission delay in the midhaul link
$D^{b,t}$ Transmission delay in the backhaul link
$R$ Data rate of the packet
$\mu$ Mean packet size
TABLE I: Notations with definitions for the system model parameters.
@@ IV Methodology
Our research addresses the resource allocation problem within the O-RAN architecture by introducing a semi-supervised learning algorithm. To achieve this, we have compiled a comprehensive dataset that includes various resource allocation scenarios, covering different objective functions and constraint combinations. We have created labeled data (computed $\alpha$, $\beta$, and transmission powers) using exhaustive search algorithm (ESA), thoroughly exploring the feasible solution space to find the global optimum solution as input for our algorithm’s supervised learning part and as a benchmark for models evaluation.
We use a deep neural network (DNN) model with a contrastive loss function to solve the resource allocation problem. This DNN model effectively manages the complexities of resource allocation and acquires the ability to map input features to resource allocation decisions. By training our model, we fine-tune it to estimate resource allocation decisions that fulfill the requirements of the optimization problem (QoS). This semi-supervised learning algorithm, which combines a VAE with a contrastive loss, is denoted as ”SS-VAE.”
We aim to evaluate the performance of our SS-VAE method using multiple regression-based performance metrics to ensure its robustness. We will also validate the effectiveness of our model by comparing its results with those of the ESA across various scenarios. While ESA is an algorithm for resource allocation decisions that find global optimum solutions, its complexity makes it impractical for real-world deployment. Therefore, we will also compare its performance with state-of-the-art methods known for their practical deployment capabilities, such as DRL. This thorough evaluation will demonstrate the reliability of our method across different scenarios.
Figure 2 illustrates our proposed methodology for training the SS-VAE and comparing its results with those of the benchmark algorithms.
Figure 2: Overview of the proposed methodology.
@@ IV-A Exhaustive Search Algorithm
The ESA is a fundamental approach used to find the global optimal solution to resource allocation problems by evaluating all possible configurations and selecting the best one based on a predefined objective function [33]. The ESA is a valuable benchmark solution that provides a ground truth against comparing results from other methods. Its advantages include simplicity of understanding and implementation and the assurance of yielding an optimal solution within a finite time frame. However, its guarantee of optimality comes at the cost of high computational effort, making it suitable primarily for small or well-defined problems where the solution space can be thoroughly explored.
Nonetheless, ESA’s main limitation lies in its computational complexity, especially when addressing resource allocation problems with high dimensions and complex constraints. The computational cost increases exponentially as the problem size grows, making ESA impractical for large-scale resource allocation tasks [34]. Despite this limitation, ESA is still widely used as a benchmark for comparing different optimization algorithms and as a reference for evaluating new optimization techniques.
To evaluate the performance, accuracy, and reliability of the SS-VAE method, we will compare its results with those obtained by ESA. We will use the ESA to obtain a subset of solutions for the objective variables ($\alpha$, $\beta$, $p$) in the resource allocation optimization problem with network slicing. This labeled dataset will be used to train the supervised learning part of our SS-VAE method. The detailed steps of ESA are described in Algorithm 1.
Algorithm 1 ESA for Resource Allocation in O-RAN
1:$B$, $U$, $S$, $M$, $w_{s}$ , $P^{max}_{b}$, $P^{max}_{s}$, $R^{min}_{s}$, $C_{b}^{max}$, $D_{u,s}^{min}$, and $H$
2:Optimal resource allocation $best\_solution$ and its objective value $best\_value$
3:Initialize dataset $\mathcal{X}\leftarrow\emptyset$
4:$best\_value\leftarrow-\infty$
5:$best\_solution\leftarrow\emptyset$
6:for each possible combination of $(\alpha,\beta,p)$ do
7:     Calculate $R_{u,s}$ using Eq. (III)
8:     if all constraints are satisfied then
9:         $current\_value\leftarrow\sum_{s\in S}\sum_{u\in U}w_{s}R_{u,s}$
10:         if $current\_value>best\_value$ then
11:              $best\_value\leftarrow current\_value$
12:              $best\_solution\leftarrow(\alpha,\beta,p)$
13:              Add $(\alpha,\beta,p,R_{u,s})$ to dataset $\mathcal{X}$
14:         end if
15:     end if
16:end for
17:return $best\_solution$, $best\_value$, $\mathcal{X}$
@@ IV-B Deep Reinforcement learning Algorithm
As discussed in Section II, DRL algorithms effectively address resource allocation, proving their practicality in wireless networks and enhancing key metrics [35]. In this study, we use DRL, where an agent allocates resources based on rewards linked to the objective in Problem 10. Hence, we model the problem as a Markov Decision Process (MDP) and use a DRL algorithm as a benchmark for evaluating our SS-VAE method. In the resource allocation problem (Equation 10), an O-RAN orchestrator is responsible for determining the PRB allocation decisions and identifying the corresponding RU for each UE within each slice. This problem can be represented as a MDP [35, 36, 37, 38].
@@
    State: The state at time $t$ ( $\mathfrak{s}(t)\in\mathfrak{S}$), is represented by $\{\mathfrak{s}_{u}(t)\}_{t=1}^{N}$, where $\mathfrak{s}_{u}(t)$ indicates the state of UE $u$ ($u=1:U$). Here, $\mathfrak{s}_{u}(t)$ is a binary indicator: it equals one if the data rate requirement of UE $u$ is satisfied and 0 otherwise. Each data rate also corresponds to a quantized transmission power level.
@@
    Reward: The reward function evaluates the gains or costs of a given state-action pair. It guides the agent’s decision-making process by indicating the desirability of various actions across different states. Consistent with previous studies [35, 36, 37], our approach defines the reward function as a combination of the objective function and the constraints pertinent to the considered problem (Equation 10). Thus, we specify the reward function for our specific scenario as follows:
@@
    where, $\Theta_{r}$, $\Theta_{const.}$, and $\Theta_{bias}$ represent the respective weights assigned to the objective function (Data Rate), constraints, and the bias value. The primary objective of the agent is to determine the optimal policy, denoted as $\pi:\mathfrak{S}\rightarrow\mathfrak{P(A)}$, for the MDP, that maximizes the cumulative reward obtained through dynamic learning from acquired data.
@@
The corresponding value function $V(\pi):\mathfrak{S}\rightarrow\mathfrak{R}$ is defined as the cumulative discounted reward achieved by following the actions prescribed by the policy $\pi$ when starting from a specific initial state. Exploring the dynamics of the environment, especially when accounting for variable channel conditions and calculating transition probabilities, is a highly complex task. Determining the value function in such scenarios presents a significant challenge. Two primary strategies are employed to address MDPs: model-based and model-free approaches. In uncertainty surrounding transition probabilities and reward functions, the model-free approach of Q-learning proves invaluable for solving the MDP [38].
Our resource allocation problem involves managing high-dimensional and complex state and action spaces due to continuous variables like transmission power and numerous UEs and PRB resources. Traditional Q-learning methods struggle in this context because of inefficient learning with rarely encountered state-action pairs [35, 38]. To solve the problem of finding parameters such as transmission power ($p$), $\alpha$, and $\beta$, we use the Deep Q-Network (DQN) approach, which handles complex state-action spaces by approximating the $Q$ function with a deep neural network. DQN benefits from experience replay and target networks, enhancing learning stability and efficiency [38]. Experience replay stores transition in memory ($\mathbb{M}$) and samples them to train the neural network ($\mathfrak{S_{t}},\mathfrak{A_{t}},\mathfrak{R_{t}},\mathfrak{S_{t+1}}$), reducing correlation in the data and improving gradient estimation. A target network stabilizes learning by updating the $Q$-network’s parameters with independent samples [38].
@@
It is worth noting that the complete training algorithm for DQN is detailed in [38], and for our specific problem, we have utilized Algorithm 3 for training the DQN model.
IV-C Proposed Method (SS-VAE)
Traditional resource allocation methods in wireless communication systems, such as heuristic and meta-heuristic algorithms, often rely on predefined rules or optimization algorithms that do not adapt well to changing network conditions. In recent times, DL techniques have been used to improve the accuracy and efficiency of resource allocation and transmission power control in wireless communication systems [39, 40, 41, 42]. These approaches are practical in learning complex, non-linear relationships between input features (e.g., channel matrices) and output decisions (e.g., power and resource allocation).
The distinctive feature of DL-based resource allocation is its capacity to manage wireless communication systems’ dynamic and complex nature effectively. DL-based approaches can adjust to changes in network conditions, user demands, and interference patterns, providing more flexible and efficient resource management than traditional methods. Traditional methods often face challenges in real-time adaptation due to predefined rules, complexity in optimization, and a lack of real-time learning. Our SS-VAE method inputs the channel gain matrix into a DNN. This network learns optimal resource allocation decisions that satisfy network constraints and maximize the weighted aggregated data rate. We use a flattened vector, with dimensions $UB$, derived from the channel gain matrix as the input for the machine learning model.
Our main contribution involves integrating a VAE with contrastive loss to solve the resource allocation problem. By adding contrastive loss into the VAE‘s loss function, we enhance the encoder’s ability to learn more refined representations, significantly improving the accuracy of resource allocation decisions. The SS-VAE method allows the encoder layers of the VAE to be explicitly utilized for the resource allocation task.
In our problem, the connection between the xApps that handle power transmissions and PRB allocation emerges from their shared foundation within the VAE framework. The VAE operates as a shared space where both xApps come together, enabling a collaborative exploration and consideration of the relevant features for both power transmission and PRB allocations.
The xApps access a common latent representation of the channel gain matrices through this shared space. This shared understanding facilitates a cohesive analysis where the relationships between power transmissions and PRB allocations are thoroughly examined, allowing for the simultaneous consideration of the key features that impact both domains. This integration of information helps make more comprehensive and informed decisions about resource allocation, as it ensures that both xApps use a unified understanding of the underlying data to optimize their respective objectives. This coordination allows for a synchronized approach to resource allocation within the O-RAN architecture.
@@ IV-C1 Variational Autoencoder
The VAE is a type of DNN generative model used in unsupervised learning tasks. It is designed to generate new data samples that resemble a provided set of training data. One notable advantage of using a VAE is its ability to adeptly capture the intrinsic distribution of the channel gain matrix (input data), which is valuable when addressing the resource allocation problem [43].
In the VAE part of our method, the input data (flattened channel gain matrix) is represented by $x$, while a lower-dimensional data representation (latent space) is denoted as $z$. The generative model based on the latent variable is represented as $p_{\theta}(x)$. The main objective of a VAE is to maximize the likelihood of the data given the model parameters, denoted as $\theta$. This likelihood can be expressed mathematically as:
@@
where $q_{\phi}(z|x)$ corresponds to the encoder network, which models the posterior distribution of the latent variable based on the input data, the term $D_{KL}(q_{\phi}(z|x)|p(z))$ represents the Kullback-Leibler (KL) divergence between this posterior distribution and the prior distribution of the latent variable. The loss function (Equation 23) can be optimized by adjusting the parameters of the generative model ($\theta$) and the encoder network represented ($\phi$). This optimization process involves iteratively updating these parameters and computing the gradient of the objective with respect to them [43]. The architecture of the VAE for addressing the resource allocation problem with network slicing is illustrated Figure 3. The VAE comprises an encoder and a decoder. The encoder compresses the channel gain matrix into a lower-dimensional latent code while the decoder reconstructs the channel gain matrix from this code. During training, the VAE minimizes reconstruction loss, improving the accuracy of predicted transmission power.
Figure 3: Architecture of the VAE.
Several key advantages can be observed when using the VAE for resource allocation. The model can represent complex distributions, capture the non-linear relationship between channel gains and resource allocation decisions, and provide a compressed latent space helpful for visualization, clustering, and anomaly detection tasks. Moreover, resource allocation problems are inherently non-convex and challenging for traditional optimization methods, where the VAEs can propose a more versatile and practical solution to addressing these complexities.
@@
The dataset used to train the SS-VAE model consists of $Q$ training samples, divided into two categories: $Q_{L}$ labeled samples and $Q_{U}$ unlabeled samples. These samples are represented by pairs of channel matrices and their corresponding power allocation vectors, denoted as $H_{i}$ and $\Gamma_{i}$ for labeled samples (where $i$ ranges from 1 to $Q_{L}$), and $H_{j}$ for unlabeled samples (where $j$ ranges from 1 to $Q_{U}$). The labeled dataset results from the ESA applied to the same training dataset. Specifically, we have used power transmission and allocation decisions ($p,\alpha,\beta$) obtained by ESA as labels for training our SS-VAE model.
@@
To train the supervised learning component of our algorithm, we employ the ADAM optimizer [44]. It facilitates the end-to-end training of the VAE using the $Q_{L}$ labeled samples. The loss function for the regression problem with ADAM optimization is formulated using the mean squared error (MSE) loss function, defined as follows:
$L_{supervised}=\frac{1}{\beta_{L}}\sum_{i\in\beta_{L}}||encoder(VAE(H_{i}))-\Gamma_{i}||^{2}.$ (24)
The encoder part of the VAE is used to estimate the transmission power and predict the allocation indicators ($\alpha_{u,b,s}$ and $\beta_{u,b,m,s}$). Our model takes a flattened vector of dimension $UB$ as input, obtained from the channel gain matrix.